OpenStudy (anonymous):

give eg.. 1. An unbounded sequence that has a convergent subsequence 2. An unbounded sequence that has no convergent subsequence 3. A null sequence (an) such that the series ¡Æ an does not converge 4. A sequence which is not a Cauchy sequence but has the property that for every ¦Å > 0 and every N > 0 there exists n > N and m > 2n such that mod(an ¨C am) ¡Ü ¦Å 5. Two sequences (an) and (bn) such that the sequence (cn) defined by cn = an + bn converges to 1 but neither (an) nor (bn) converge. 6. A sequence (an) that tends to +ve infinity but is neither increasing nor eventually inc

OpenStudy (anonymous):

1. (-1)^n

OpenStudy (anonymous):

\((-1)^n\) is a bounded sequence... You can use a different kind of alternating sequence that's unbounded, but still contains a convergent subsequence, such as \[\left\{1,\pi,2,\pi,3,\pi,...\right\}\] The particular subsequence of interest is \(\left\{\pi,\pi,\pi,...\right\}\), which obviously converges to \(\pi\).

OpenStudy (anonymous):

For (5), you can consider the following sequences: \[\{a_n\}=\{-1,1,-1,1,...\}\\ \{b_n\}=\{2,0,2,0\}\] Then you have \(a_1+b_1=a_2+b_2=a_3+b_3=\cdots=1\), so \(\{c_n\}=\{1,1,1,1,...\}\), yet neither \(a_n\) nor \(b_n\) converge.

OpenStudy (anonymous):

For (2), I would think the natural numbers \(\{1,2,3,...\}\) would suffice, but not totally sure.